OpenStudy (anonymous):

An investor invested a total of $1900 in two mutual funds. One fund earned a 4 percent profit while the other earned a 2 percent profit. If the invester's total profit was $64 how much was invested in each mutual fund?

OpenStudy (ranga):

Assume the investment in the 4% profit fund is x Then the investment in the 2% profit fund is (1900 - x) Profit from the 4% fund = (x)(4/100) = 0.04x Profit from the 2% fund = (1900 - x)(2/100) = 38 - 0.02x Add up the two profits and equate it to $64 and solve for x.